# Low-Altitude Wireless Networks: A Survey

### v1

@@ 1.1 Background
Traditional terrestrial networks provide foundational capabilities like communication, localized sensing (e.g., global navigation satellite system (GNSS), cellular triangulation), and fixed-point computation, supporting ubiquitous connectivity and basic data services [1]. However, these existing networks are proving increasingly inadequate to meet the unique demands of emerging low-altitude applications. While terrestrial networks excel at their original design goals, their inherent limitations, such as static infrastructure and limited sensing capabilities, restrict their ability to effectively support dynamic and mobile sensing, communication, computation, and control in three-dimensional airspace. This inability to adequately support low-altitude applications stems from several key limitations. First, terrestrial networks primarily offer static sensing and communication, failing to provide the dynamic, high-resolution, and mobile sensing vital for many low-altitude use cases. This makes them unsuitable for applications that require real-time environmental monitoring in diverse locations, on-demand data collection in hard-to-reach areas, or physical manipulation at varying altitudes [2]. Second, the performance of low-altitude applications often depends on balancing complex, multi-objective requirements that cannot be managed using finite-level performance metrics. Terrestrial networks face major challenges to find end-to-end optimization that simultaneously balances data timeliness, terminal energy, and network scalability [3], especially in the dynamic and often unpredictable conditions of low-altitude environments. Moreover, the open nature of wireless channels makes terrestrial networks susceptible to malicious attacks. Relying on single-dimensional security and privacy measures, they are highly vulnerable to sophisticated, multi-dimensional attacks that compromise data integrity, breach privacy, cause denial-of-service, or prevent trusted communication [4]. These vulnerabilities become particularly acute in low-altitude applications, where security is paramount for safety-critical operations.
With the rapid emergence of a low-altitude digital services ecosystem, low-altitude wireless networks (LAWNs) are increasingly recognized as essential infrastructure for building three-dimensional intelligent airspace. As illustrated in Fig. 1, LAWNs are designed to provide integrated sensing, communication, computation and control services for aircraft (drones, electric vertical take-off and landing (eVTOLs), etc.) operating in the airspace below 3000 meters above ground level [5]. The core mission of LAWNs is to restructure the airspace operation paradigm through wireless technology, thereby supporting the safe and efficient operation of the low-altitude economy [6], alleviating the strain on terrestrial network development, and offering technical insights for their further evolution. The detailed insights are listed as follows:
@@
  • LAWNs achieve a synergistic integration of sensing, computing, and control within a network of mobile unmanned aerial vehicles (UAVs) [7]. This distributed approach allows for on-demand, high-resolution, and mobile environmental perception, going beyond the capabilities of static terrestrial sensors. Furthermore, it enables payload delivery or infrastructure inspection, supporting a wider range of low-altitude use cases. This distributed and dynamic sensing and interaction surpasses the limitations of static and restricted sensing/interaction characteristics of terrestrial networks, providing crucial support for low-altitude applications.
@@
  • LAWNs excel in delivering diverse services to UAVs, offering far more flexible configurations than terrestrial networks can provide [5]. As illustrated in Fig. 1, they dynamically adapt parameters such as UAV altitude, speed, trajectory, communication settings, and computational offloading strategies. This enables the simultaneous enhancement of data timeliness, energy efficiency, network coverage, communication reliability, and operational safety. This adaptable, system-level performance is often unattainable with the statically optimized configurations of terrestrial infrastructure, and unlocks new capabilities for aerial vehicles.
@@
  • Security is a paramount concern for LAWNs, as safe and reliable operation is a prerequisite to enable a robust low-altitude economy. To this end, LAWNs leverage advanced attack defense and privacy preservation strategies, capitalizing on their mobile and distributed architecture to enable dynamic deployment of security mechanisms [8]. These include adaptive jamming resistance, secure multi-drone collaboration for threat detection, cutting-edge encryption protocols, and privacy-preserving data processing techniques. This adaptable approach yields robust, multi-dimensional security and privacy solutions specifically tailored to the unique vulnerabilities and operational contexts of aerial platforms, ensuring the safe and reliable operation of low-altitude services.
@@ 1.2 Motivation
There have been several surveys exploring the LAWN designs from various perspectives in terms of resource allocation, energy efficiency, and learning-based techniques. Taking an example, the studies in [5, 25, 20] highlighted the flexibility and on-demand connectivity offered by LAWNs, while emphasizing the advancing capabilities of security, robust sensing, communication, and computing. The authors in [18] introduced LAWN to enhance edge intelligence by supporting edge computing and caching services simultaneously, offering solutions to service provisioning challenges in 6G networks. Regarding techniques for learning-based LAWN designs, the works of [26, 11] comprehensively examined the application of artificial intelligence (AI) to autonomous multi-UAV wireless networks (MUWN). Furthermore, [15] focused on energy harvesting (EH) techniques for LAWNs, elaborating how EH can extend LAWN sustainability and facilitate connectivity for energy-constrained devices, particularly in remote or disaster-stricken areas. Recent surveys resorted to exploring the interplay between LAWN and key technologies for 6G. For instance, the study in [16] detailed the integration of reconfigurable intelligent surfaces (RIS) to enhance LAWN system performance. With the ability to manipulate signal transmission, RIS-based LAWN can improve UAV limitations in areas like coverage extension, signal reliability, and energy efficiency. The authors in [17] investigated the integrated sensing and communication (ISAC) design LAWN, demonstrating the superiority in inference management and target tracking. Beyond communication-centric research, efforts have also been devoted to multiple complementary dimensions of LAWNs. For instance, the study in [9] evaluated various control methods of LAWNs in dynamic and uncertain environments, including model predictive control (MPC) and proportional-integrative-derivative (PID). Besides, the study in [23] discussed the fundamental hardware components of multi-copter UAVs, outlining key methods for optimizing critical parameters to ensure reliable and effective UAV usage. The work in [10] further summarized convex formulations for air traffic flow optimization, presenting a solution to addressing challenges such as congestion and capacity constraints.
@@
For ease of presentation, we compare the related surveys in the TABLE 1. It can be seen that on the one hand, the rise of LAWNs can bring plenty of advantages such as improved coverage in underserved areas, flexible deployment options, and enhanced line-of-sight communications. On the other hand, due to the challenges of high-volume, multi-connection data, and the open nature of wireless channels, key concerns in terms of reliable communications, accurate sensing, and robust control need further joint consideration. It is, however, noted that the related surveys do not fully consider the multiple functions, various indicators, and intertwined privacy and security issues of the LAWNs holistically. Moreover, the practical implementation of LAWNs requires an investigation about the low-altitude airspace (LAA) structuring and management. In this regard, existing literature with limited scopes of LAWNs may not provide a complete understanding of the trade-offs between performance, functionality, and security/privacy, and potentially result in suboptimal designs and deployments for real-world LAWN applications. To close up the above research gap, this article provides an overview of LAWN from functionality, performance evaluation, as well as air traffic management (ATM), offering a more comprehensive and broader perspective on the design, deployment, and management of LAWN systems. The main contributions of the paper can be summarized as follows:
@@
As outlined in Fig. 2, the rest of the paper is organized as follows. Section 2 introduces the fundamentals of LAWNs, including LAWN architecture and LAWN channel modeling. Then, Section 3 focuses on the multi-functions of the LAWNs, mainly from sensing, computing, communication, control, and energy-delivering aspects. Section 4 sheds light on the multi-indicators of the LAWNs from timely, energy, scalability, and connectivity perspectives. Subsequently, Section 5 illustrates the privacy preservation, physical layer security, and network layer security issues of LAWN. Section 6 talks about the LAWN implementation and management. Finally, Sections 7 and 8 present the future research challenges and briefly conclude the paper, respectively.
@@ 2.1 LAWN Architecture
1) Aerial platform: The heterogeneous low-altitude platforms, including both rotary/fixed-wing drones and uncrewed/manned aircraft, serve as the most important role of LAWNs, performing tasks such as wireless control, distributed sensing. Based on flight endurance and operational radius, aerial platforms can be simply classified into long-endurance UAVs (e.g., CH-4, Global Hawk) for strategic surveillance, with coverage exceeding 1000 km, mid-range UAVs for tactical reconnaissance and environmental monitoring, covering several hundred kilometers; and micro and consumer-grade UAVs (e.g., DJI Mavic series) with limited flight durations and control via remote consoles or mobile apps, typically operating within 5-30 km.
2) Terrestrial Infrastructure: The terrestrial infrastructure forms the backbone to support practical LAWN realizations. In particular, the takeoff and landing places like vertiports are equipped with necessary turnaround supports for battery charging/refueling and repairing. Meanwhile, the ground base stations (GBS) are responsible for data transmission conveying mission planning. In addition, the environmental services, including geofencing, high-definition (HD) maps, and weather forecasts, provide crucial inputs for predictive designs.
@@
It is evident that aerial-to-ground (A2G) communications are indispensable functions within LAWNs. However, we herein highlight that while LAWN systems are established on UAV-based platforms, they differ fundamentally from conventional UAV communication architectures. Therefore, it is critical to clarify the scope of LAWNs to avoid confusion. Functionally, traditional UAV communication systems emphasize wireless transmission performance at the physical layer, where aerial nodes typically act as terminal users or relays within pre-established cellular or ad hoc networks, with minimal focus on flight attitude control, battery management, and formation coordination. In contrast, LAWNs extend beyond traditional communication purposes by supporting the seamless integration of sensing, computation, and control, enabling real-time coordination across multiple domains. It is not intended to replace dedicated aerial communication services, which may still be offloaded to conventional networks. Instead, LAWN focuses on the co-design of networking models, control-aware protocols, spatial scheduling, and real-time decision-making across airspace-aware nodes. Moreover, it emphasizes large-scale network coordination, enabling collaborative efforts across multiple airspace-aware nodes to optimize performance and efficiency [27]. Furthermore, LAWNs require comprehensive optimization beyond physical layers. For instance, efficient routing and network protocols must be co-designed to accommodate dynamic drone mobility and task-driven communication needs, while application layer designs ensure that mission-critical tasks can be real-time controlled by remote apps.
Figure 3: LAWN architectures including star networks with direct A2G links, multi-group mesh networks with intra-group cooperation, and hierarchical mesh networks with multi-layer A2A connections for scalable coverage.
@@
To ensure the effective functioning of LAWNs, it is essential to establish seamless and efficient information flow, both between the ground crew and UAVs and amongst the UAVs themselves. This, in turn, demands a robust and adaptive network architecture that can support not only high-capacity data transfer but also real-time coordination of various operational elements. Consequently, we proceed to focus on LAWN architectures ranging from star to mesh networks shown in Fig. 3, particularly emphasizing multi-group and hierarchical mesh networks [20, 28].
@@
2) Mesh Networks: As the scope and complexity of LAWNs continue to expand, the inherent limitations of star network topologies become increasingly evident. To overcome these limitations, mesh networks provide a decentralized communication architecture. In a mesh network, each LAWN node can communicate directly with other nodes via air-to-air (A2A) links, either via direct communication or through multiple hops. The decentralized structure ensures that even if one node fails or moves out of range, the network can adapt and maintain connectivity. As a consequence, mesh networks are particularly well-suited for LAWN systems to maintain robustness and reliability in high-mobility environments. Ad Hoc networks naturally form as a subset of mesh networks, where each LAWN node autonomously communicates with nearby nodes, establishing a peer-to-peer communication model. As discussed in [29], the flexibility of Ad Hoc networks ensures that LAWN nodes can adapt to changes in topology and network conditions, enabling continuous communication even in remote areas where communication infrastructure is unavailable. While a single mesh network offers significant improvements over star networks, the multi-group mesh network architecture provides an effective solution for larger LAWN systems. In a multi-group mesh network, multiple independent mesh networks operate in parallel, each serving a subset of LAWN nodes. Communication between these groups is facilitated by backbone nodes that connect the different mesh networks, ensuring that data can flow seamlessly across the entire LAWNs. Moreover, compared to a single mesh network, multi-group networks allow for efficient load balancing, reducing congestion and improving overall network performance.
@@
For even larger systems, hierarchical mesh networks offer an efficient solution [19]. In a hierarchical mesh network, nodes are organized into several hierarchical layers, each responsible for different communication tasks. The lower layers consist of local mesh networks, where LAWN nodes communicate within their group. Higher layers, composed of more powerful nodes, manage inter-group communication and connect the network to the ground station or other higher-tier nodes. This hierarchical structure enhances the network’s scalability and fault tolerance by distributing the communication load across different layers. The multi-layer design also improves network efficiency by optimizing communication paths, reducing congestion at the higher layers, and ensuring that the network remains operational even if one or more layers fail. The hierarchical approach is particularly beneficial for missions involving large numbers of LAWN nodes with varying communication capabilities, such as self-organized drone formation working in coordination for search and rescue operations or wide-area surveillance.

### v2

# Low-Altitude Wireless Networks: A Comprehensive Survey

@@ 1.1 Background
Traditional terrestrial networks provide foundational capabilities like communication, localized sensing (e.g., global navigation satellite system (GNSS) and cellular triangulation), and fixed-point computation, supporting ubiquitous connectivity and basic data services [asadi2014survey]. However, these existing networks are proving increasingly inadequate to meet the unique demands of emerging low-altitude applications. While terrestrial networks excel at their original design goals, their inherent limitations, such as static infrastructure and limited sensing capabilities, restrict their ability to effectively support dynamic and mobile sensing, communication, computation, and control in three-dimensional airspace. This inability to adequately support low-altitude applications stems from several key limitations. First, terrestrial networks primarily offer static sensing and communication, failing to provide the dynamic, high-resolution, and mobile sensing vital for many low-altitude use cases. This makes them unsuitable for applications that require real-time environmental monitoring in diverse locations, on-demand data collection in hard-to-reach areas, or physical manipulation at varying altitudes [wang2019convergence]. Second, the performance of low-altitude applications often depends on balancing complex, multi-objective requirements that cannot be managed using finite-level performance metrics. Terrestrial networks face major challenges to find end-to-end optimization that simultaneously balances data timeliness, terminal energy, and network scalability [wang2025bridging], especially in the dynamic and often unpredictable conditions of low-altitude environments. Moreover, the open nature of wireless channels makes terrestrial networks susceptible to malicious attacks. Relying on single-dimensional security and privacy measures, they are highly vulnerable to sophisticated, multi-dimensional attacks that compromise data integrity, breach privacy, cause denial-of-service, or prevent trusted communication [mahmood2019analysis]. These vulnerabilities become particularly acute in low-altitude applications, where security is paramount for safety-critical operations.
With the rapid emergence of a low-altitude digital services ecosystem, low-altitude wireless networks (LAWNs) are increasingly recognized as essential infrastructure for building three-dimensional intelligent airspace. As illustrated in Fig. 1, LAWNs are designed to provide integrated sensing, communication, computation and control services for aircraft (drones, electric vertical take-off and landing (eVTOLs), etc.) operating in the airspace below 3000 meters above ground level [wang2025toward]. The core mission of LAWNs is to restructure the airspace operation paradigm through wireless technology, thereby supporting the safe and efficient operation of the low-altitude economy (LAE) [liao2024benefits], alleviating the strain on terrestrial network development, and offering technical insights for their further evolution. The detailed insights are listed as follows:
@@
  • LAWNs achieve a synergistic integration of sensing, computing, and control within a network of mobile unmanned aerial vehicles (UAVs) [yang2024embodied]. This distributed approach allows for on-demand, high-resolution, and mobile environmental perception, going beyond the capabilities of static terrestrial sensors. Furthermore, it enables payload delivery or infrastructure inspection, supporting a wider range of low-altitude use cases. This distributed and dynamic sensing and interaction surpasses the limitations of static and restricted sensing/interaction characteristics of terrestrial networks, providing crucial support for low-altitude applications.
@@
  • LAWNs excel in delivering diverse services to UAVs, offering far more flexible configurations than terrestrial networks can provide [wang2025toward]. As illustrated in Fig. 1, they dynamically adapt parameters such as UAV altitude, speed, trajectory, communication settings, and computational offloading strategies. This enables the simultaneous enhancement of data timeliness, energy efficiency, network coverage, communication reliability, and operational safety. This adaptable, system-level performance is often unattainable with the statically optimized configurations of terrestrial infrastructure, and unlocks new capabilities for aerial vehicles.
@@
  • Security is a paramount concern for LAWNs, as safe and reliable operation is a prerequisite to enable a robust low-altitude economy. To this end, LAWNs leverage advanced attack defense and privacy preservation strategies, capitalizing on their mobile and distributed architecture to enable dynamic deployment of security mechanisms [cai2025secure]. These include adaptive jamming resistance, secure multi-drone collaboration for threat detection, cutting-edge encryption protocols, and privacy-preserving data processing techniques. This adaptable approach yields robust, multi-dimensional security and privacy solutions specifically tailored to the unique vulnerabilities and operational contexts of aerial platforms, ensuring the safe and reliable operation of low-altitude services.
@@
Table 2: List of Acronyms
Acronym Description
5G Fifth-generation
6G Sixth-generation
A2A Air-to-air
A2G Air-to-ground
ADS-B Automatic dependent surveillance–broadcast
AI Artificial intelligence
ATM Air traffic management
ATC Air traffic control
CNPC Control and non-payload communication
D2D Device-to-device
EH Energy harvesting
eVTOL Electric vertical take-off and landing
GBS Ground base station
GNSS Global navigation satellite system
HAP High-altitude platform
ISAC Integrated sensing and communication
IoT Internet of Things
LAE Low-altitude economy
LAA Low-altitude airspace
LAWN Low-altitude wireless network
LiDAR Light detection and ranging
LLM Large language model
LoS Line-of-sight
LTE Long-term evolution
MARL Multi-agent reinforcement learning
MIMO Multiple-input multiple-output
mmWave Millimeter wave
MEC Mobile edge computing
MPC Model predictive control
MFD Macroscopic fundamental diagram
NLoS Non-line-of-sight
NOMA Non-orthogonal multiple access
PID Proportional–integral–derivative
RF Radio frequency
RIS Reconfigurable intelligent surface
UAV Unmanned aerial vehicle
UAM Urban air mobility
WPT Wireless power transfer
@@ 1.2 Motivation
There have been several surveys exploring the LAWN designs from various perspectives in terms of resource allocation, energy efficiency, and learning-based techniques. Taking an example, the studies in [wang2025toward, dai2022unmanned, 119] highlighted the flexibility and on-demand connectivity offered by LAWNs, while emphasizing the advancing capabilities of security, robust sensing, communication, and computing. The authors in [9520337] introduced LAWN to enhance edge intelligence by supporting edge computing and caching services simultaneously, offering solutions to service provisioning challenges in 6G networks. Regarding techniques for learning-based LAWN designs, the works of [bai2023toward, kurunathan2023machine] comprehensively examined the application of artificial intelligence (AI) to autonomous multi-UAV wireless networks (MUWN). Furthermore, [pandey2024uav] focused on energy harvesting (EH) techniques for LAWNs, elaborating how EH can extend LAWN sustainability and facilitate connectivity for energy-constrained devices, particularly in remote or disaster-stricken areas. Recent surveys resorted to exploring the interplay between LAWN and key technologies for 6G. For instance, the study in [pogaku2022uav] detailed the integration of reconfigurable intelligent surfaces (RIS) to enhance LAWN system performance. With the ability to manipulate signal transmission, RIS-based LAWN can improve UAV limitations in areas like coverage extension, signal reliability, and energy efficiency. The authors in [meng2023uav] investigated the integrated sensing and communication (ISAC) design LAWN, demonstrating the superiority in inference management and target tracking. Beyond communication-centric research, efforts have also been devoted to multiple complementary dimensions of LAWNs. For instance, the study in [jin2025advancing] evaluated various control methods of LAWNs in dynamic and uncertain environments, including model predictive control (MPC) and proportional-integrative-derivative (PID). Besides, the study in [sdoukou2025survey] discussed the fundamental hardware components of multi-copter UAVs, outlining key methods for optimizing critical parameters to ensure reliable and effective UAV usage. The work in [Work2009] further summarized convex formulations for air traffic flow optimization, presenting a solution to addressing challenges such as congestion and capacity constraints.
@@
For ease of presentation, we compare the related surveys in TABLE 1. It can be seen that on the one hand, the rise of LAWNs can bring plenty of advantages such as improved coverage in underserved areas, flexible deployment options, and enhanced line-of-sight communications. On the other hand, due to the challenges of high-volume, multi-connection data, and the open nature of wireless channels, key concerns in terms of reliable communications, accurate sensing, and robust control need further joint consideration. It is, however, noted that the related surveys do not fully consider the multiple functions, various indicators, and intertwined privacy and security issues of the LAWNs holistically. Moreover, the practical implementation of LAWNs requires an investigation about the low-altitude airspace (LAA) structuring and management. In this regard, existing literature with limited scopes of LAWNs may not provide a complete understanding of the trade-offs between performance, functionality, and security/privacy, and potentially result in suboptimal designs and deployments for real-world LAWN applications. To close up the above research gap, this article provides an overview of LAWN from functionality and performance evaluation, as well as air traffic management (ATM), offering a more comprehensive and broader perspective on the design, deployment, and management of LAWN systems. The main contributions of the paper can be summarized as follows:
@@
The remainder of the paper is structured as shown in Fig. 2. Section 2 introduces the fundamentals of LAWNs, including system architecture and representative channel models. Then, section 3 discusses LAWN performance metrics in terms of timeliness and latency, energy-efficient design, scalability, and connectivity. In section 4, we present multi-functional LAWN design methods. Subsequently, section 5 is devoted to multi-level defenses of LAWNs, covering privacy preservation, physical-layer security, and network-layer security. Section 6 focuses on advanced AI-driven LAWNs, including both large-model-based and agentic AI-based architectures. Section 7 addresses implementation and management aspects in terms of airspace structuring, traffic management, and path planning. Finally, the open research challenges and conclusions are outlined in sections 8 and 9, respectively. For brevity, we further summarize the acronyms of this survey in TABLE 2.
@@ 2.1 LAWN Architecture
1) Aerial platform: The aerial platforms form the physical backbone of LAWNs, providing the carriers for wireless control, distributed sensing, and service delivery. Following the classification in [atkins2009commercial, dalamagkidis2008unmanned], these platforms can be distinguished along several complementary dimensions. In terms of flight altitude, operations typically span low-, medium-, and high-altitude regimes within the sub-3000 m airspace. From a weight perspective, platforms range from micro to large aircraft and include both lighter-than-air vehicles and heavier-than-air fixed/rotor-wing designs. Mission profile and deployment area further differentiate platforms used in rural, suburban, and dense urban environments for tasks such as surveillance, entertainment, humanitarian relief, and swarm operations. Finally, range and endurance give rise to long-endurance systems capable of flights beyond 1000 km, mid-range platforms covering several hundred kilometers, and consumer-grade drones that typically operate within about 5–30 km under remote or app-based control. To sum up, the classification of drones under various criteria is demonstrated in Fig. 3.
Figure 3: Classification of drones under various criteria.
2) Terrestrial Infrastructure: The terrestrial infrastructure forms the backbone to support practical LAWN realizations. In particular, the takeoff and landing places like vertiports are equipped with necessary turnaround supports for battery charging/refueling and repairing. Meanwhile, the ground base stations (GBS) are responsible for data transmission conveying mission planning. In addition, the environmental services, including geofencing, high-definition maps, and weather forecasts, provide crucial inputs for predictive designs.
@@
It is evident that aerial-to-ground (A2G) communications are indispensable functions within LAWNs. However, we herein highlight that while LAWN systems are established on UAV-based platforms, they differ fundamentally from conventional UAV communication architectures. Therefore, it is critical to clarify the scope of LAWNs to avoid confusion. Functionally, traditional UAV communication systems emphasize wireless transmission performance at the physical layer, where aerial nodes typically act as terminal users or relays within pre-established cellular or Ad-hoc networks, with minimal focus on flight attitude control, battery management, and formation coordination. In contrast, LAWNs extend beyond traditional communication purposes by supporting the seamless integration of sensing, computation, and control, enabling real-time coordination across multiple domains. It is not intended to replace dedicated aerial communication services, which may still be offloaded to conventional networks. Instead, LAWN focuses on the co-design of networking models, control-aware protocols, spatial scheduling, and real-time decision-making across airspace-aware nodes. Moreover, it emphasizes large-scale network coordination, enabling collaborative efforts across multiple airspace-aware nodes to optimize performance and efficiency [yuan2025ground]. Furthermore, LAWNs require comprehensive optimization beyond physical layers. For instance, efficient routing and network protocols must be co-designed to accommodate dynamic drone mobility and task-driven communication needs, while application layer designs ensure that mission-critical tasks can be real-time controlled by remote apps.
From a broader networking perspective, LAWNs are also conceptually distinct from space-air-ground integrated networks (SAGINs), which are designed to provide global coverage by vertically integrating satellite, aerial, and terrestrial segments into a unified communication infrastructure [chen2024survey]. SAGIN research primarily concentrates on multi-layer connectivity, backhaul routing, and resource sharing across heterogeneous layers, while low-altitude airspace management, mission-centric control, and interaction with urban environments are only indirectly addressed. LAWNs, on the other hand, focus on the sub-3000 m domain and are tailored for dense, task-driven operations that tightly couple airspace structuring, traffic management, sensing, computation, and control. Rather than being regarded as a subset of SAGIN, LAWNs are better interpreted as a complementary low-altitude operational layer that can interconnect with SAGIN for backbone connectivity, while maintaining control-aware protocols and airspace-aware service for LAE applications.
Table 3: Comparison of representative LAWN architectures
Topology Reliability Latency Resource efficiency Scalability
Star network Vulnerable to central-node failure Low for nearby nodes, grows with hub load and distance High for small-scale networks with simple centralized routing Limited; performance degrades as the node number increases
Mesh network Moderate; enabling local rerouting under failures Moderate due to multi-hop forwarding Moderate; incurring additional signaling and energy cost Good for medium-size deployments
Multi-group mesh High; backbone links connecting groups Low for intra-group; moderate for inter-group Improved efficiency through balanced load across groups High; partitioning into groups supports large systems with manageable overhead
Hierarchical mesh High; tiered redundancy and fault isolation Low; high-tier links offer short end-to-end paths for wide-area communication High; aggregated control and traffic management enhance global resource usage Very high; suitable for large and heterogeneous LAWNs with diverse service demands
@@
To ensure the effective functioning of LAWNs, it is essential to establish seamless and efficient information flow, both between the ground crew and UAVs and amongst the UAVs themselves. This, in turn, demands a robust and adaptive network architecture that can support not only high-capacity data transfer but also real-time coordination of various operational elements. Consequently, we proceed to focus on LAWN architectures ranging from star to mesh networks shown in Fig. 4, particularly emphasizing multi-group and hierarchical mesh networks [119, akyildiz2005wireless].
@@
2) Mesh Networks: As the scope and complexity of LAWNs continue to expand, the inherent limitations of star network topologies become increasingly evident. To overcome these limitations, mesh networks provide a decentralized communication architecture. In a mesh network, each LAWN node can communicate directly with other nodes via air-to-air (A2A) links, either via direct communication or through multiple hops. The decentralized structure ensures that even if one node fails or moves out of range, the network can adapt and maintain connectivity. As a consequence, mesh networks are particularly well-suited for LAWN systems to maintain robustness and reliability in high-mobility environments. Ad-hoc networks naturally form as a subset of mesh networks, where each LAWN node autonomously communicates with nearby nodes, establishing a peer-to-peer communication model. As discussed in [120], the flexibility of Ad-hoc networks ensures that LAWN nodes can adapt to changes in topology and network conditions, enabling continuous communication even in remote areas where communication infrastructure is unavailable. While a single mesh network offers significant improvements over star networks, the multi-group mesh network architecture provides an effective solution for larger LAWN systems. In a multi-group mesh network, multiple independent mesh networks operate in parallel, each serving a subset of LAWN nodes. Communication between these groups is facilitated by backbone nodes that connect the different mesh networks, ensuring that data can flow seamlessly across the entire LAWNs. Moreover, compared to a single mesh network, multi-group networks allow for efficient load balancing, reducing congestion and improving overall network performance.
Figure 4: LAWN architectures, including star networks with direct A2G links, multi-group mesh networks with intra-group cooperation, and hierarchical mesh networks with multi-layer A2A connections for scalable coverage.
@@
For even larger systems, hierarchical mesh networks offer an efficient solution [127]. In a hierarchical mesh network, nodes are organized into several hierarchical layers, each responsible for different communication tasks. The lower layers consist of local mesh networks, where LAWN nodes communicate within their group. Higher layers, composed of more powerful nodes, manage inter-group communication and connect the network to the ground station or other higher-tier nodes. This hierarchical structure enhances the network’s scalability and fault tolerance by distributing the communication load across different layers. The multi-layer design also improves network efficiency by optimizing communication paths, reducing congestion at the higher layers, and ensuring that the network remains operational even if one or more layers fail. The hierarchical approach is particularly beneficial for missions involving large numbers of LAWN nodes with varying communication capabilities, such as self-organized drone formations working in coordination for search and rescue operations or wide-area surveillance. To sum up, the representative LAWN architectures discussed above are compared in TABLE 3 in terms of reliability, latency, resource efficiency, and scalability.